OpenStudy (anonymous):

Find the value of the annuity \[a_1 = 9000, I=0.05, n= 10\]

OpenStudy (anonymous):

This is what I have so far \[9000 (\frac{(1+0.05)^{10 }-1 }{ 0.05 })\] I think that this would be the right approach

OpenStudy (amistre64):

future value? or present value?

OpenStudy (anonymous):

future

OpenStudy (amistre64):

looks fine to me

OpenStudy (anonymous):

If i do math on the top I got 0.62889463 do I divide that with the bottom 0.05

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

I got 12.5778926

OpenStudy (amistre64):

that is what would be called a factor of 1 dollar.

OpenStudy (anonymous):

do I then multiply this number to the 9000

OpenStudy (amistre64):

yes, since there is 9000 dollars, and the factor of 1 dollar is 12.5779 multiply the 2 to get the value for 9000 dollars

OpenStudy (anonymous):

I got 113201.1

OpenStudy (anonymous):

so $113,201.10

OpenStudy (amistre64):

seems reasonable to me

OpenStudy (amistre64):

i got .03 or .04, but thats immaterial

OpenStudy (anonymous):

It says in my homework that this isn't right

OpenStudy (anonymous):

I work in mathlab so if the answer is wrong it tells me

OpenStudy (amistre64):

it also has a "work similar problem" option. make sure your assumptions are correct, or if its wanting you to use a tabled value for the factor instead

OpenStudy (amistre64):

it might even be saying to round to the nearest yadayada

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

round to the nearest cent

OpenStudy (amistre64):

does it ask for a tabled factor, or a formula factor? if no table is presented, then we can assume the formula method

OpenStudy (anonymous):

there is no table

OpenStudy (amistre64):

y1 = P(1+r)^0 y2 = P(1+r)^1 + P(1+r)^0 y3 = P(1+r)^2 + P(1+r)^1 + P(1+r)^0 .... yn = P(1+r)^(n-1))+ ... +P(1+r)^2 + P(1+r)^1 + P(1+r)^0 let 1+r = k for cleaning up yn = P(k)^(n-1))+ ... + P(k)^2 + P(k)^1 + P(k)^0 yn = P(k^(n-1)+ ... +(k)^2 + (k)^1 + (k)^0) \[y_n = P\frac{1-k^n}{1-k}\] \[y_n = P\frac{1-(1+.05)^n}{1-(1+.05)}\] \[y_n = P\frac{1-(1+.05)^n}{-.05}\] \[y_n = P\frac{(1+.05)^n-1}{.05}\]

OpenStudy (amistre64):

9000((1.05)^10-1)/.05 = 113 201 . 032 82

OpenStudy (anonymous):

You have me lost sorry

OpenStudy (amistre64):

i just went and worked out the formula .... then i inputed the specifics

OpenStudy (anonymous):

oh ok